# Outlier Gradient Analysis: Efficiently Identifying Detrimental Training Samples for Deep Learning Models

### v3

@@ 9 Conclusion
In this paper, we focused on the key data-centric learning task of identifying detrimental training samples. Influence functions are a leading approach often used for this problem, but possess certain deficiencies when applied to deep models. As part of our proposed solution, we address the key challenge associated with influence functions usage in deep models— the computational demands for inverting the Hessian matrix. This results in a computationally efficient method that possesses high detection accuracy. Our approach, outlier gradient analysis, is based on a conceptual transformation between the influence function formulation and outlier analysis in the gradient space. Through comprehensive experiments on synthetic datasets and various application domains (code details in Appendix E), including noisy label correction for vision models, data selection for NLP models, and even influential data identification in LLMs, we demonstrated that our method outperformed many existing influence-based approaches and baselines in deep learning scenarios.
@@ Appendix A Additional Related Work on Miscellaneous Data-Centric Learning
Many works in the data-centric learning domain study other relevant research questions beyond detrimental sample identification and influence estimation. For instance, datamodels [Ilyas et al., 2022] also estimate training sample contributions, but only for one test sample at a time. Data efficiency approaches [Jain et al., 2023, Paul et al., 2021, Killamsetty et al., 2021] aim to accelerate deep learning training time via subset selection. Data pruning approaches based on novel approximations for leave-one-out influence estimation [Tan et al., 2024] and the model’s generalization gap [Yang et al., 2022] have also been proposed. Model pruning via generalized influence functions has also been studied in Lyu et al. [2023]. Note that after identifying detrimental training samples, one can adopt multiple strategies for recourse. While we focus on removal in this paper, other alternatives could also be used, such as relabeling [Richardson et al., 2023, Kong et al., 2021]. Antidote data augmentation [Chhabra et al., 2022, Li et al., 2023] methods aim to generate synthetic data samples to improve model performance, whereas feature selection approaches [Hall, 1999, Cai et al., 2018] seek to optimize the feature space to only those important for model performance. Active learning [Cohn et al., 1996] methods aim to iteratively identify optimal samples to annotate given a large unlabeled training data pool [Liu et al., 2021, Nguyen et al., 2022, Wei et al., 2015]. Finally, works on poisoning attacks seek to analyze model robustness by perturbing training set samples [Solans et al., 2021, Mehrabi et al., 2021, Chhabra et al., 2023] under natural input constraints. The study of training sample influence has also been extended to recent generative models, such as diffusion models [Dai and Gifford, 2023], through the use of ensembles.
@@ B.1.1 Synthetic Datasets
We conduct experiments for our proposed outlier gradient analysis and other baselines on two synthetic datasets. The first dataset is linearly separable for logistic regression classification and consists of 150 training samples and 100 test samples. These are created using the scikit-learn [Pedregosa et al., 2011] library’s make_blobs function. For each of the two binary classes, we manually flip the labels of 10 samples (5 for each class) to add noise to the dataset. The second dataset is the non-linear half moons dataset so that we can train an MLP network with two hidden layers with ReLU activations. The training set has 250 samples and the test set has 100 samples, and the dataset is generated using the scikit-learn library’s make_moons function. Here too, we manually flip the labels of 20 samples (10 from each class) to add noise to the data.
@@ B.1.2 CIFAR-10N and CIFAR-100N
Both the CIFAR-10N and CIFAR-100N datasets [Wei et al., 2022] consist of the same input images that make up the CIFAR-10 (10 classes) and CIFAR-100 (100 classes) datasets [Krizhevsky et al., 2009], respectively. Each input is a 32x32 RGB image with dimension (3,32,32). However, for CIFAR-10N and CIFAR-100N, the labels are noisy, as they contain real-world human annotation errors collected using 3 annotators on Amazon Mechanical Turk. As these datasets are based on human-annotated noise, they model noisy real-world datasets more realistically, compared to synthetic data alternatives. The training set for both datasets contains 50,000 image-label pairs, and the test set contains 10,000 image-label pairs that are free from noise. For CIFAR-10N we utilize three noise settings for experiments in the paper– (1) Worst, which is the dataset version with the highest noise rate (40.21%) as the worst possible annotation label for the image is chosen, (2) Aggregate, which is the least noisy dataset (9.03%) as labels are chosen via majority voting amongst the annotations, and (3) Random which has intermediate noise (17.23%) and consists of picking one of the annotators’ labels. We use the first annotator for the random labels. For CIFAR-100N there is only a single noisy setting (Noisy100) due to the large number of labeling classes, and the overall noise rate is 40.20%.
@@ B.1.3 GLUE Datasets
The GLUE or the General Language Understanding Evaluation [Wang et al., 2018] benchmark datasets consist of a number of benchmarks for training, evaluating, and analyzing natural language models. As in the DataInf paper [Kwon et al., 2024], we utilize the four binary classification subset datasets: QNLI, SST2, QQP, and MRPC for experiments. Here, these datasets cover a wide variety of natural language task domains. For instance, QNLI [Wang et al., 2018] covers natural language inference, SST2 [Socher et al., 2013] covers sentiment analysis, QQP  [^4] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs covers question answering, and MRPC [Dolan and Brockett, 2005] covers paraphrase detection. We use the same datasets as in Kwon et al. [2024], where the training and test splits are obtained from the Huggingface datasets [^5] https://huggingface.co/docs/datasets library. For QQP and SST2 in Kwon et al. [2024] 4500 training samples and 500 test samples were randomly sampled from the full sets, so we utilize these in our experiments for a fair comparison.
@@ B.1.4 Sentence Transformations
For this benchmark dataset proposed in Kwon et al. [2024], the LLM is required to perform a specific transformation on an input sentence. There are 10 different sentence transformations. To help the model learn different transformations, “chatbot” name identifiers are used and each is uniquely associated with each transformation. These are the categories of sentence transformations (taking an example input sentence as “Welcome to the real world.”):
@@ B.2 Models and Methods
We now describe the models and the methods used in our experiments throughout the main paper. First, we describe the ResNet-34 [He et al., 2016] architecture used as the base model for the noisy vision datasets, then the RoBERTa [Liu et al., 2019] NLP transformer model, and then the Llama-2 LLM. [^6] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf. We also describe implementation details and parameter values for the label correction baselines in Sections 4 and 5 and the influence-based baselines used throughout the paper. Finally, we also describe some key implementation details regarding our outlier gradient analysis approach.
@@ B.2.1 ResNet-34
The ResNet-34 model was proposed in He et al. [2016] and is a 34-layer convolutional neural network pretrained on the ImageNet-1K dataset at resolution 224 $\times$ 224. The pretrained model block is fine-tuned on the CIFAR-10N/CIFAR-100N training set experiments with default parameters– minibatch size (128), optimizer (SGD), initial learning rate (0.1), momentum (0.9), weight decay (0.0005), and number of epochs (100), for all experiments. Moreover, we directly used the implementation provided by Wei et al. [2022] and made modifications to their code.
@@ B.2.2 RoBERTa
As in Kwon et al. [2024], we utilize LoRA fine-tuning to fine-tune the RoBERTa-large model, a 355M parameter transformer language model that improves upon the original BERT model in key ways such as implementation and hyperparameter selection. LoRA is applied to every value matrix of the attention layers of the RoBERTa model. The pre-trained model from Huggingface is used. [^7] https://huggingface.co/docs/transformers/model_doc/roberta. A learning rate of 0.0003 and a batch size of 32 is used. The model is fine-tuned over 10 epochs using LoRA and dropout is set to be 0.05 while the rank of the LoRA matrix is set to 4, as recommended in Kwon et al. [2024]. The loss function used is a negative log-likelihood as the datasets are all for binary classification. The LoRA training is enabled using the Huggingface PEFT library. [^8] https://huggingface.co/docs/peft/index. For the influence experiments we have utilized the code provided in Kwon et al. [2024] and adapted it for our experiments. Moreover, we only compute influences using the training set gradients, and keep the test set hidden from the learning model for fair evaluation.
@@ B.2.3 Llama2-13B-chat LLM
We fine-tune the Llama2 13B parameter instruction tuned LLM using LoRA fine-tuning (applied to every query and value matrix of the attention layer) as in Kwon et al. [2024]. The LoRA parameters are as follows: learning rate is set to be 0.0003, rank of LoRA matrix is set to 8, $\alpha=32$ in 8-bit quantization, and the batch size is set to 32 across 25 fine-tuning epochs. A negative log-likelihood of the generated response is used as the loss function for fine-tuning as before. Here too, we adapt the code provided by Kwon et al. [2024] for our use cases.
@@ B.2.4 Label Correction Baselines
For label correction baselines in Sections 4 and 5Normalized Margin [Northcutt et al., 2021], Self-Confidence [Müller and Markert, 2019], and Confidence-Weighted Entropy [Kuan and Mueller, 2022], we utilize the implementation provided in the Cleanlab [^9] https://github.com/cleanlab/cleanlab/. library. We use default parameters for all three baselines. Note that the baselines are model agnostic and only require predicted labels and associated probabilities for predictions, which we can easily obtain from classifiers.
@@ B.2.5 Influence-Based Baselines
We utilize three influence-based baselines in experiments: LiSSA [Koh and Liang, 2017], Gradient Tracing [Pruthi et al., 2020], DataInf [Kwon et al., 2024]. For each of these baselines, we utilize the implementation provided in Kwon et al. [2024] and adapt it to our application scenarios. For each baseline influence estimation is undertaken only on the training set.
@@ B.2.6 Outlier Gradient Analysis
We now discuss implementation details regarding outlier gradient analysis. Owing to the simplicity of our approach, the implementation is straightforward and follows directly from the algorithm. In most cases, we directly utilize the gradients obtained from the last layer of the model being considered. However, in some cases, the gradient space of samples can be high dimensional. For instance, for CIFAR-100N, the gradient space is of dimension 50000 $\times$ 51200 which unnecessarily increases memory and time complexity of outlier detection. As a result, we reduce the gradient space dimensionality by employing a sparse random projection step [Li et al., 2006] where the reduced dimension is ascertained using the scikit-learn library. We also utilize sparse random projection in this manner for the Llama-2-13B-chat LLM model experiments to reduce the dimensionality of the gradient space obtained.
@@ Appendix C Additional Results and Experiments
Normalized Margin [Northcutt et al., 2021] 91.33 ± 0.11 90.06 ± 0.14 83.57 ± 0.32 60.94 ± 0.59
Self-Confidence [Müller and Markert, 2019] 91.38 ± 0.19 90.09 ± 0.17 83.65 ± 0.21 60.51 ± 0.51
Confidence Entropy [Kuan and Mueller, 2022] 91.11 ± 0.34 90.05 ± 0.26 83.63 ± 0.41 60.62 ± 0.26
Gradient Tracing [Pruthi et al., 2020] 91.47 ± 0.21 89.98 ± 0.20 83.38 ± 0.58 60.73 ± 0.38
LiSSA [Koh and Liang, 2017] 91.49 ± 0.34 90.05 ± 0.31 83.38 ± 0.58 60.48 ± 0.29
DataInf [Kwon et al., 2024] 91.46 ± 0.17 90.05 ± 0.38 83.40 ± 0.56 60.70 ± 0.31
Self-LiSSA [Bejan et al., 2023] 92.07 ± 0.15 89.58 ± 0.11 83.01 ± 0.34 59.48 ± 0.43
@@ C.3 Experiments on Running Time and Computational Complexity
LiSSA [Koh and Liang, 2017] Hessian-based $\mathcal{O}(nvp)$
DataInf [Kwon et al., 2024] Hessian-based $\mathcal{O}(nvp)$
EK-FAC LLM Baseline [Grosse et al., 2023] Hessian-based $\mathcal{O}(m^{2}o+p^{2}o)$
Self-LiSSA [Bejan et al., 2023] Self-influence $\mathcal{O}(np)$
Self-DataInf Self-influence $\mathcal{O}(np)$
Gradient Tracing Hessian-free $\mathcal{O}(nvp)$

### v5

@@ 9 Conclusion
We focused on the key data-centric learning task of identifying detrimental training samples. Influence functions are a leading approach often used for this problem, but possess certain deficiencies when applied to deep models, such as the computational demands for inverting the Hessian matrix. We propose a novel solution for detrimental sample detection that does not rely on the Hessian matrix, and hence eliminates this major limitation. Our approach, outlier gradient analysis, is based on a conceptual transformation between the influence function formulation and outlier analysis in the gradient space. This transformation results in a computationally efficient method that possesses high detection accuracy. Through comprehensive experiments on synthetic datasets and various application domains (code details in Appendix D), including noisy label correction for vision models, data selection for NLP models, and even influential data identification in LLMs, we demonstrated that our method outperformed many existing influence-based approaches and baselines in deep learning scenarios.
10 Broader Impact and Limitations
Our work and proposed techniques aim to address the data-centric task of identifying detrimental samples. We improve upon the influence function analysis framework that is used to undertake this problem, but possesses deficiencies when applied to deep learning models. Enabling influence estimation for deep models allows practitioners to assess whether training samples are beneficial or detrimental to performance, and can make models more interpretable and performant. As we show through extensive experiments on multiple problem settings, our proposed outlier gradient analysis approach outperforms existing baselines and can augment model performance by identifying/trimming detrimental samples in a computationally efficient manner. As a result, our work paves the way for significant positive societal impact, especially with the increased adoption of larger and deeper neural networks such as LLMs. However, as with any work, there are limitations to our approaches that can be overcome in future work. For instance, it might be possible to derive specific outlier analysis algorithms that are computationally more efficient than iForest or norm thresholding, and significantly more performant. Another limitation that can be overcome is the further study and benchmarks for influence based analysis in LLMs– going beyond the datasets and approaches we used in this work. Further, while outlier gradient analysis is useful in cases where training data can be noisy, it might not be as useful if the data is very high quality and there are no outlying gradient samples. However, it is unlikely that this will be the case in the real-world unless some steps have been taken prior to training to ensure high data quality. Finally, outlier analysis algorithms have a fundamental limitation of how to specify the budget for outlier detection, which is a non-trivial hyperparameter optimization problem. While this is a common problem with little consensus across the entire field of outlier analysis, our methods inherit this limitation as well (although we note that outlier gradient analysis performs well for different budget thresholds, as shown in additional experiments in the Appendix C.2).
Acknowledgments
The authors would like to thank Han Yue for aiding with experiment design and implementation, and the anonymous reviewers for their feedback in helping strengthen the work. Anshuman Chhabra was supported by the USF CSE department faculty startup fund for the duration of this project.
@@ Appendix A Additional Related Work on Miscellaneous Data-Centric Learning
Many works in the data-centric learning domain study other relevant research questions beyond detrimental sample identification and influence estimation. For instance, datamodels (Ilyas et al., 2022) also estimate training sample contributions, but only for one test sample at a time. Data efficiency approaches (Jain et al., 2023; Paul et al., 2021; Killamsetty et al., 2021) aim to accelerate deep learning training time via subset selection. Data pruning approaches based on novel approximations for leave-one-out influence estimation (Tan et al., 2024) and the model’s generalization gap (Yang et al., 2022) have also been proposed. Model pruning via generalized influence functions has also been studied in (Lyu et al., 2023). Note that after identifying detrimental training samples, one can adopt multiple strategies for recourse. While we focus on removal in this paper, other alternatives could also be used, such as relabeling (Richardson et al., 2023; Kong et al., 2021). Antidote data augmentation (Chhabra et al., 2022; Li et al., 2023) methods aim to generate synthetic data samples to improve model performance, whereas feature selection approaches (Hall, 1999; Cai et al., 2018) seek to optimize the feature space to only those important for model performance. Active learning (Cohn et al., 1996) methods aim to iteratively identify optimal samples to annotate given a large unlabeled training data pool (Liu et al., 2021; Nguyen et al., 2022; Wei et al., 2015). Finally, works on poisoning attacks seek to analyze model robustness by perturbing training set samples (Solans et al., 2021; Mehrabi et al., 2021; Chhabra et al., 2023) under natural input constraints. The study of training sample influence has also been extended to recent generative models, such as diffusion models (Dai & Gifford, 2023), through the use of ensembles.
@@ B.1.1 Synthetic Datasets
We conduct experiments for our proposed outlier gradient analysis and other baselines on two synthetic datasets. The first dataset is linearly separable for logistic regression classification and consists of 150 training samples and 100 test samples. These are created using the scikit-learn (Pedregosa et al., 2011) library’s make_blobs function. For each of the two binary classes, we manually flip the labels of 10 samples (5 for each class) to add noise to the dataset. The second dataset is the non-linear half moons dataset so that we can train an MLP network with two hidden layers with ReLU activations. The training set has 250 samples and the test set has 100 samples, and the dataset is generated using the scikit-learn library’s make_moons function. Here too, we manually flip the labels of 20 samples (10 from each class) to add noise to the data.
@@ B.1.2 CIFAR-10N and CIFAR-100N
Both the CIFAR-10N and CIFAR-100N datasets (Wei et al., 2022) consist of the same input images that make up the CIFAR-10 (10 classes) and CIFAR-100 (100 classes) datasets (Krizhevsky et al., 2009), respectively. Each input is a 32x32 RGB image with dimension (3,32,32). However, for CIFAR-10N and CIFAR-100N, the labels are noisy, as they contain real-world human annotation errors collected using 3 annotators on Amazon Mechanical Turk. As these datasets are based on human-annotated noise, they model noisy real-world datasets more realistically, compared to synthetic data alternatives. The training set for both datasets contains 50,000 image-label pairs, and the test set contains 10,000 image-label pairs that are free from noise. For CIFAR-10N we utilize three noise settings for experiments in the paper– (1) Worst, which is the dataset version with the highest noise rate (40.21%) as the worst possible annotation label for the image is chosen, (2) Aggregate, which is the least noisy dataset (9.03%) as labels are chosen via majority voting amongst the annotations, and (3) Random which has intermediate noise (17.23%) and consists of picking one of the annotators’ labels. We use the first annotator for the random labels. For CIFAR-100N there is only a single noisy setting (Noisy100) due to the large number of labeling classes, and the overall noise rate is 40.20%.
@@ B.1.3 GLUE Datasets
The GLUE or the General Language Understanding Evaluation (Wang et al., 2018) benchmark datasets consist of a number of benchmarks for training, evaluating, and analyzing natural language models. As in the DataInf paper (Kwon et al., 2024), we utilize the four binary classification subset datasets: QNLI, SST2, QQP, and MRPC for experiments. Here, these datasets cover a wide variety of natural language task domains. For instance, QNLI (Wang et al., 2018) covers natural language inference, SST2 (Socher et al., 2013) covers sentiment analysis, QQP  [^2] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs covers question answering, and MRPC (Dolan & Brockett, 2005) covers paraphrase detection. We use the same datasets as in Kwon et al. (2024), where the training and test splits are obtained from the Huggingface datasets [^3] https://huggingface.co/docs/datasets library. For QQP and SST2 in Kwon et al. (2024) 4500 training samples and 500 test samples were randomly sampled from the full sets, so we utilize these in our experiments for a fair comparison.
@@ B.1.4 Sentence Transformations
For this benchmark dataset proposed in (Kwon et al., 2024), the LLM is required to perform a specific transformation on an input sentence. There are 10 different sentence transformations. To help the model learn different transformations, “chatbot” name identifiers are used and each is uniquely associated with each transformation. These are the categories of sentence transformations (taking an example input sentence as “Welcome to the real world.”):
@@ B.2 Models and Methods
We now describe the models and the methods used in our experiments throughout the main paper. First, we describe the ResNet-34 (He et al., 2016) architecture used as the base model for the noisy vision datasets, then the RoBERTa (Liu et al., 2019) NLP transformer model, and then the Llama-2 LLM. [^4] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf. We also describe implementation details and parameter values for the label correction baselines in Sections 4 and 5 and the influence-based baselines used throughout the paper. Finally, we also describe some key implementation details regarding our outlier gradient analysis approach.
@@ B.2.1 ResNet-34
The ResNet-34 model was proposed in (He et al., 2016) and is a 34-layer convolutional neural network pretrained on the ImageNet-1K dataset at resolution 224 $\times$ 224. The pretrained model block is fine-tuned on the CIFAR-10N/CIFAR-100N training set experiments with default parameters– minibatch size (128), optimizer (SGD), initial learning rate (0.1), momentum (0.9), weight decay (0.0005), and number of epochs (100), for all experiments. Moreover, we directly used the implementation provided by Wei et al. (2022) and made modifications to their code.
@@ B.2.2 RoBERTa
As in (Kwon et al., 2024), we utilize LoRA fine-tuning to fine-tune the RoBERTa-large model, a 355M parameter transformer language model that improves upon the original BERT model in key ways such as implementation and hyperparameter selection. LoRA is applied to every value matrix of the attention layers of the RoBERTa model. The pre-trained model from Huggingface is used. [^5] https://huggingface.co/docs/transformers/model_doc/roberta. A learning rate of 0.0003 and a batch size of 32 is used. The model is fine-tuned over 10 epochs using LoRA and dropout is set to be 0.05 while the rank of the LoRA matrix is set to 4, as recommended in Kwon et al. (2024). The loss function used is a negative log-likelihood as the datasets are all for binary classification. The LoRA training is enabled using the Huggingface PEFT library. [^6] https://huggingface.co/docs/peft/index. For the influence experiments we have utilized the code provided in (Kwon et al., 2024) and adapted it for our experiments. Moreover, we only compute influences using the training set gradients, and keep the test set hidden from the learning model for fair evaluation.
@@ B.2.3 Llama2-13B-chat LLM
We fine-tune the Llama2 13B parameter instruction tuned LLM using LoRA fine-tuning (applied to every query and value matrix of the attention layer) as in Kwon et al. (2024). The LoRA parameters are as follows: learning rate is set to be 0.0003, rank of LoRA matrix is set to 8, $\alpha=32$ in 8-bit quantization, and the batch size is set to 32 across 25 fine-tuning epochs. A negative log-likelihood of the generated response is used as the loss function for fine-tuning as before. Here too, we adapt the code provided by Kwon et al. (2024) for our use cases.
@@ B.2.4 Label Correction Baselines
For label correction baselines in Sections 4 and 5Normalized Margin (Northcutt et al., 2021), Self-Confidence (Müller & Markert, 2019), and Confidence-Weighted Entropy (Kuan & Mueller, 2022), we utilize the implementation provided in the Cleanlab [^7] https://github.com/cleanlab/cleanlab/. library. We use default parameters for all three baselines. Note that the baselines are model agnostic and only require predicted labels and associated probabilities for predictions, which we can easily obtain from classifiers.
@@ B.2.5 Influence-Based Baselines
We utilize three influence-based baselines in experiments: LiSSA (Koh & Liang, 2017), Gradient Tracing (Pruthi et al., 2020), DataInf (Kwon et al., 2024). For each of these baselines, we utilize the implementation provided in Kwon et al. (2024) and adapt it to our application scenarios. For each baseline influence estimation is undertaken only on the training set (except for additional results in adapting to the test set, provided in Appendix C.9 below). We only use the last checkpoint in Gradient Tracing (Pruthi et al., 2020) for fair comparisons.
@@ B.2.6 Outlier Gradient Analysis
We now discuss implementation details regarding outlier gradient analysis. Owing to the simplicity of our approach, the implementation is straightforward and follows directly from the algorithm. In most cases, we directly utilize the gradients obtained from the last layer of the model being considered. However, in some cases, the gradient space of samples can be high dimensional. For instance, for CIFAR-100N, the gradient space is of dimension 50000 $\times$ 51200 which unnecessarily increases memory and time complexity of outlier detection. As a result, we reduce the gradient space dimensionality by employing a sparse random projection step (Li et al., 2006) where the reduced dimension is ascertained using the scikit-learn library. We also utilize sparse random projection in this manner for the Llama-2-13B-chat LLM model experiments to reduce the dimensionality of the gradient space obtained.
@@ Appendix C Additional Results and Experiments
Normalized Margin (Northcutt et al., 2021) 91.33 ± 0.11 90.06 ± 0.14 83.57 ± 0.32 60.94 ± 0.59
Self-Confidence (Müller & Markert, 2019) 91.38 ± 0.19 90.09 ± 0.17 83.65 ± 0.21 60.51 ± 0.51
Confidence Entropy (Kuan & Mueller, 2022) 91.11 ± 0.34 90.05 ± 0.26 83.63 ± 0.41 60.62 ± 0.26
Gradient Tracing (Pruthi et al., 2020) 91.47 ± 0.21 89.98 ± 0.20 83.38 ± 0.58 60.73 ± 0.38
LiSSA (Koh & Liang, 2017) 91.49 ± 0.34 90.05 ± 0.31 83.38 ± 0.58 60.48 ± 0.29
DataInf (Kwon et al., 2024) 91.46 ± 0.17 90.05 ± 0.38 83.40 ± 0.56 60.70 ± 0.31
Self-LiSSA (Bejan et al., 2023) 92.07 ± 0.15 89.58 ± 0.11 83.01 ± 0.34 59.48 ± 0.43
@@ C.3 Experiments on Running Time and Computational Complexity
LiSSA (Koh & Liang, 2017) Hessian-based $\mathcal{O}(nvp)$
DataInf (Kwon et al., 2024) Hessian-based $\mathcal{O}(nvp)$
EK-FAC (Grosse et al., 2023) Hessian-based $\mathcal{O}(m^{2}o+p^{2}o)$
Self-LiSSA (Bejan et al., 2023) Self-influence $\mathcal{O}(np)$
Self-DataInf (Bejan et al., 2023) Self-influence $\mathcal{O}(np)$
Gradient Tracing (Pruthi et al., 2020), Hessian-free $\mathcal{O}(nvp)$